OpenStudy (anonymous):

f(x) = x3 + 5x2 - x - 5 Part A: What are the factors of f(x)? Show your work. (3 points)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@SolomonZelman @Preetha

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf f(x) = x^3 + 5x^2 - x - 5 }\) \(\Large\color{red}{ \bf f(x) = x^3 - x + 5x^2- 5 }\) \(\Large\color{red}{ \bf f(x) = x(x^2 - 1) + 5(x^2- 1) }\) \(\Large\color{red}{ \bf f(x) = (x+5)(x^2 - 1) }\) \(\Large\color{red}{ \bf f(x) = (x+5)(x - 1)(x+1) }\)

OpenStudy (solomonzelman):

you can do the rest.

OpenStudy (anonymous):

sothe factors would be f(x)=(x+5)(x−1)(x+1)

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf anytime! }\)